OpenStudy (anonymous):

the sum of two consecutive page numbers in a book is 223 find the page numbers

OpenStudy (gw2011):

Let x=the higher page number Then x-1=the lower page number x+(x-1)=223 x+x-1=223 2x-1=223 2x=223+1 2x=224 x=112 x-1=111 Therefore, the page numbers are 111 and 112